OpenStudy (anonymous):

448 students and 32 teachers to 234 students and 18 teachers witch is the best ratio

OpenStudy (mathstudent55):

What is better, many students per teacher or few students per teacher?

OpenStudy (anonymous):

many students per teacher

OpenStudy (mathstudent55):

I would say that most teachers, parents, and students would consider a low number of students per teacher to be better. This way the teacher has more time to dedicate to each student.

OpenStudy (mathstudent55):

Let's find in each case what the ratio is of students per teacher. 448 students and 32 teachers is a ratio of \(\dfrac{448}{32} \) which reduces to \(\dfrac{14}{1}\) or 14 students per teacher. 234 students and 18 teachers is a ratio of \(\dfrac{234}{18} \) which reduces to \(\dfrac{13}{1}\) or 13 students per teacher.

OpenStudy (mathstudent55):

Now you pick which is the better ratio.

OpenStudy (anonymous):

OK thank you for helping me c:

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

hello are you still there